# DAWN-FM: Data-Aware and Noise-Informed Flow Matching for Solving Inverse Problems

### v2

@@ Abstract
Inverse problems, which involve estimating parameters from incomplete or noisy observations, arise in various fields such as medical imaging, geophysics, and signal processing. These problems are often ill-posed, requiring regularization techniques to stabilize the solution. In this work, we employ Flow Matching (FM), a generative framework that integrates a deterministic processes to map a simple reference distribution, such as a Gaussian, to the target distribution. Our method DAWN-FM: Data-AWare and Noise-informed Flow Matching incorporates data and noise embedding, allowing the model to access representations about the measured data explicitly and also account for noise in the observations, making it particularly robust in scenarios where data is noisy or incomplete. By learning a time-dependent velocity field, FM not only provides accurate solutions but also enables uncertainty quantification by generating multiple plausible outcomes. Unlike pre-trained diffusion models, which may struggle in highly ill-posed settings, our approach is trained specifically for each inverse problem and adapts to varying noise levels. We validate the effectiveness and robustness of our method through extensive numerical experiments on tasks such as image deblurring and tomography.
1 Introduction
Figure 1: Example images from MNIST (left) and STL10 (right) datasets showing the recovery of deblurred images from the blurred (noisy) data (top panel) at different levels of noise (0%, 1%, 5%, 10%, 20%). Our methods DAW-FM  (middle panel) incorporates the embedding for the blurred data within the network, while DAWN-FM (bottom panel) incorporates embedding for both data and noise levels. DAWN-FM is superior to DAW-FM at deblurred image recovery, especially at higher levels of noise.
@@ 1 Introduction
Inverse problems are a class of problems in which the goal is to determine parameters (or parameter function) of a system from observed data. These problems arise in various fields, including medical imaging, geophysics, remote sensing, and signal processing. Inverse problems are often ill-posed, meaning that a unique solution does not exist, or the solution may be highly sensitive to small perturbations in the data. Solving inverse problems typically requires regularization techniques, which introduce additional constraints or prior information to stabilize the solution and mitigate the effects of ill-posedness. Such regularization can be obtained using various mathematical and computational methods, including optimization techniques (variational methods), statistical inference (Bayesian or frequentist) and machine learning [Scherzer et al., 2009, Adler and Öktem, 2017, Calvetti and Somersalo, 2024, López-Tapia et al., 2021]. In this paper, we approach inverse problems using the latter and investigate a new set of methods that are machine learning based for the solution of inverse problems. In particular, we show how flow matching (FM) which was recently proposed in the context of generative models can be effectively used to estimate the solution of inverse problems and to further investigate its non-uniqueness.
@@
Stochastic Interpolation (SI) is a relatively new generative process that provides a unifying framework, elegantly integrating both deterministic flows and stochastic diffusion models [Albergo and Vanden-Eijnden, 2022, Albergo et al., 2023]. The core concept of SI is to learn a stochastic process that effectively transports a simple reference distribution, such as Gaussian, to the desired target data distribution. This transportation process can manifest as either deterministic or stochastic. In the former case, it is described by an ordinary differential equation (ODE), while in the latter, it is governed by a stochastic differential equation (SDE). For this work, we rely only on the formal deterministic ODE formalism of SI, which is known as Flow Matching (FM).
@@
Related work: The methods proposed here are closely related to three different approaches for the solution of inverse problems. First, there is an obvious link to the incorporation of diffusion models as regularizers in inverse problems [Yang et al., 2022, Chung et al., 2022a, c, b, Song et al., 2022]. The key idea is to leverage a pre-trained diffusion model that captures the data distribution as a prior, and then condition this model on the given measurements to infer the underlying clean signal or image. For inverse problems, diffusion methods condition the diffusion model on the given measurements (e.g. noisy, incomplete, or compressed data) by incorporating them into the denoising process. Nonetheless, it has been shown in Eliasof et al. [2024] that pre-trained diffusion models that are used for ill-posed inverse problems as regularizers tend to under-perform as compared to the models that are trained specifically on a particular inverse problem. In particular, such models tend to break when the noise level is not very low.
A second branch of techniques that are related to the work proposed here use encoder-decoder type networks for the solution of inverse problems (see Chung et al. [2024], Thomas et al. [2022] and reference within). Such approaches are sometimes referred to as likelihood-free estimators as they yield a solution without the computation of the likelihood. This technique is particularly useful for problems where the forward problem is difficult to compute.
A third branch of techniques that relates to our approach uses the forward problem within the neural network (see Eliasof et al. [2024], Mukherjee et al. [2021], Eliasof et al. [2023], Jin et al. [2017] and reference within). In this approach, one computes the data misfit (and its gradient) within the network and use it to guide training. Our approach utilizes components from this methodology to deal with the measured data.
@@
Main contribution: The core contributions of this paper lie in the design and application of the DAWN-FM framework, specifically crafted for solving highly ill-posed inverse problems like image deblurring and tomography. (i) Our method is designed to be problem-specific, which adjusts itself to the unique structure of the inverse problem it is tasked with, learning the posterior distribution directly, ensuring that the learned velocity fields and mappings are directly applicable to the target task. (ii) We train an interpolant by embedding measured data and noise information directly into the interpolation process. This incorporation allows our model to adapt to a wide range of noise conditions, something that pre-trained models struggle with. By making the training explicitly aware of the noise level and data characteristics, our model can better navigate noisy or incomplete measured data, producing superior reconstructions. (iii) By learning the posterior distribution directly and leveraging the stochastic nature of the interpolation process, DAWN-FM generates multiple plausible solutions for a given inverse problem, allowing us to explore the solution space more thoroughly and in particular, to estimate the posterior mean and its standard deviation, estimating the uncertainty in the recovered solution.
2 Flow Matching and Inverse Problems
@@
2.1 Flow Matching (FM): A partial review
Flow Matching (FM) is a framework that transforms points between two distributions. Given two densities $\pi_{0}({\bf x})$ and $\pi_{1}({\bf x})$, the goal is to find a mapping that takes a point ${\bf x}_{0}\sim\pi_{0}({\bf x})$ and transports it to a point ${\bf x}_{1}\sim\pi_{1}({\bf x})$. For simplicity and for the purpose of this work, we choose $\pi_{0}$ to be a Gaussian distribution with 00 mean and ${\bf I}$ covariance.
Consider sampling points from both distributions and define the trajectories
${\bf x}_{t}=(1-t){\bf x}_{0}+t{\bf x}_{1}$ (1)
These trajectories connect points from ${\bf x}_{0}$ at $t=0$ to ${\bf x}_{1}$ at $t=1$. More complex trajectories have been proposed in Albergo et al. [2023], however, in our context, we found that simple linear trajectories suffice and have advantages for being very smooth in time. The velocity along the trajectory is the time derivative of the trajectory, that is,
${\frac{d{\bf x}_{t}}{dt}}={\bf v}={\bf x}_{1}-{\bf x}_{0}$ (2)
Under the FM framework, the velocity field for all $({\bf x}_{t},t)$ is learned by averaging over all possible paths. To this end, we parameterize the velocity by a function ${\bf s}_{{\boldsymbol{\theta}}}({\bf x}_{t},t)$ and solve the optimization problem for ${\boldsymbol{\theta}}$,
$\widehat{\boldsymbol{\theta}}={\rm arg}\min_{{\boldsymbol{\theta}}}{\mathbb{E}}_{{\bf x}_{0},{\bf x}_{1}}\Big{[}\|{\bf s}_{\theta}({\bf x}_{t},t)-{\bf v}\|^{2}\Big{]}={\rm arg}\min_{{\boldsymbol{\theta}}}{\mathbb{E}}_{{\bf x}_{0},{\bf x}_{1}}\Big{[}\|{\bf s}_{\theta}({\bf x}_{t},t)+{\bf x}_{0}-{\bf x}_{1}\|^{2}\Big{]}$ (3)
After training, the velocity function ${\bf s}_{{\boldsymbol{\theta}}}({\bf x}_{t},t)$ can be estimated for every point in time. In this work, given ${\bf x}_{0}$, we recover ${\bf x}_{1}$ by numerically integrating the ODE
@@ 2.1 Flow Matching (FM): A partial review
${\bf x}_{1}={\bf x}_{t}+(1-t){\bf v}\approx{\bf x}_{t}+(1-t){\bf s}_{{\boldsymbol{\theta}}}({\bf x}_{t},t)$ (5)
@@
2.2 Applying FM for solving inverse problems
@@
${\bf A}{\bf x}_{1}+{\boldsymbol{\varepsilon}}={\bf b}$ (6)
Here, ${\bf A}$ is a linear forward mapping (although the method can work for nonlinear mappings as well) and ${\boldsymbol{\varepsilon}}\sim\mathcal{N}(0,\sigma^{2}{\bf I})$ is a random vector. We assume that ${\bf A}$ is rank-deficient or numerically rank-deficient [Hansen, 1997], so the effective dimension of ${\bf b}$ is smaller than ${\bf x}_{1}$ and one cannot obtain a reasonable estimate for ${\bf x}_{1}$ given the noisy data ${\bf b}$ without the incorporation of a-priori information.
@@
Bayes’ theorem suggests that it is possible to factor the posterior distribution $\pi({\bf x}_{1}|{\bf b})$ using the known distribution of $\pi({\bf b}|{\bf x}_{1})$ and the prior distribution $\pi({\bf x}_{1})$. This observation motivated a number of studies that used the estimated pre-trained distribution $\pi({\bf x}_{1})$ in the process of solving an inverse problem [Yang et al., 2022, Chung et al., 2022a, c, b, Song et al., 2022]. Nonetheless, it has been shown in Eliasof et al. [2024] that such estimators tend to produce unsatisfactory results when solving highly ill-posed problems or when the data is very noisy. The reason for this behaviour stems from the fact that pre-trained estimators push the solution towards the center of the prior distribution, irrespective of what the data represents. To see this, we use a careful dissection of the solution. Consider the singular value decomposition
${\bf A}=\sum_{i}{\boldsymbol{\lambda}}_{i}{\bf u}_{i}{\bf v}_{i}^{\top}$
@@
where ${\bf u}_{i}$ and ${\bf v}_{i}$ are the left and right singular vectors and ${\boldsymbol{\lambda}}_{i}$ are the singular values. Given the orthogonality of ${\bf u}$, we can decouple the data equations into
${\boldsymbol{\lambda}}_{i}({\bf v}_{i}^{\top}{\bf x}_{1})=({\bf u}_{i}^{\top}{\bf b}),\quad i=1,\ldots,n$
@@
If ${\boldsymbol{\lambda}}_{i}$ is large, the projection of ${\bf x}_{1}$ onto the eigenvector ${\bf v}_{i}$ is very informative and very minimal regularization is required. However, if ${\boldsymbol{\lambda}}_{i}\approx 0$, the contribution of ${\bf v}_{i}$ to the solution is difficult, if not impossible, to obtain and this is where the regularization is highly needed. When using pre-trained models, the prior $\pi({\bf x}_{1})$ is estimated numerically and it is unaware of the inverse problem at hand. Errors in the estimated $\pi({\bf x}_{1})$ in the parts that correspond to the large singular vectors may not be destructive. However, if $\pi({\bf x}_{1})$ has errors that correspond to the very small singular values, that is, to the effective null space of the data, this may lead to the artifacts that have been observed early in Kaipio and Somersalo [2004], Tenorio et al. [2011]. This suggests that, although it is appealing to use a generic pre-trained priors in the process, a better approach is to not use the Bayesian factorization to prior and likelihood but rather train a flow matching interpolant that maps the distribution $\pi({\bf x}_{0})$ to the posterior $\pi({\bf x}_{1}|{\bf b})$ directly. Indeed, as we show next, the flexibility of the FM framework allows us to learn a velocity function that achieves just that.
2.3 A data-aware and noise-informed velocity estimator

### v3

# arXiv:2412.04766v3[eess.IV] 28 Feb 2026

@@ Abstract.
Inverse problems, which involve estimating parameters from incomplete or noisy observations, arise in various fields such as medical imaging, geophysics, and signal processing. These problems are often ill-posed, requiring regularization techniques to stabilize the solution. In this work, we employ Flow Matching (FM), a generative framework that integrates a deterministic processes to map a simple reference distribution, such as a Gaussian, to the target distribution. Our method DAWN-FM: Data-AWare and Noise-Informed Flow Matching incorporates data and noise embedding, allowing the model to access representations about the measured data explicitly and also account for noise in the observations, making it particularly robust in scenarios where data is noisy or incomplete. By learning a time-dependent velocity field, FM not only provides accurate solutions but also enables uncertainty quantification by generating multiple plausible outcomes. Unlike pretrained diffusion models, which may struggle in highly ill-posed settings, our approach is trained specifically for each inverse problem and adapts to varying noise levels. We validate the effectiveness and robustness of our method through extensive numerical experiments on tasks such as image deblurring and tomography. The code is available at: https://github.com/ahxmeds/DAWN-FM.git.
Key words and phrases:
Inverse problems, Flow matching, Stochastic interpolation, Image deblurring, Tomography, Data embedding, Noise embedding
1991 Mathematics Subject Classification:
Primary: 65J20, 65J22, 68T07; Secondary: 65C30.
The first author is supported by Canadian Institutes of Health Research (CIHR) Grant PIBH-GR018169.
Corresponding author: Shadab Ahamed
[…]
1Department of Physics & Astronomy, University of British Columbia, Vancouver, BC Canada
2Department of Earth, Ocean and Atmospheric Sciences, University of British Columbia, Vancouver, BC Canada
(Communicated by Carola-Bibiane Schönlieb)
1. Introduction
@@
Inverse problems are a class of problems in which the goal is to determine parameters (or parameter function) of a system from observed data. These problems arise in various fields, including medical imaging, geophysics, remote sensing, and signal processing. Inverse problems are often ill-posed, meaning that a unique solution does not exist, or the solution may be highly sensitive to small perturbations in the data. Solving inverse problems typically requires regularization techniques, which introduce additional constraints or prior information to stabilize the solution and mitigate the effects of ill-posedness. Such regularization can be obtained using various mathematical and computational methods, including optimization techniques (variational methods), statistical inference (Bayesian or frequentist) and machine learning [Scherzer2009, adler2017solving, calvetti2024distributed, LOPEZTAPIA2021103285]. In this paper, we approach inverse problems using the latter and investigate a new set of methods that are machine learning based for the solution of inverse problems. In particular, we show how flow matching (FM) which was recently proposed by Lipman et al. [lipman2022flow] in the context of generative models can be effectively used to estimate the solution of inverse problems and to further investigate its non-uniqueness.
@@
Stochastic Interpolation (SI) is a relatively new generative process that provides a unifying framework, elegantly integrating both deterministic flows and stochastic diffusion models [albergo2022building, albergo2023stochastic]. The core concept of SI is to learn a stochastic process that effectively transports a simple reference distribution, such as Gaussian, to the desired target data distribution. This transportation process can manifest as either deterministic or stochastic. In the former case, it is described by an ordinary differential equation (ODE), while in the latter, it is governed by a stochastic differential equation (SDE). For this work, we rely only on the formal deterministic ODE formalism of SI, which is known as Flow Matching (FM).
@@
Related work: The methods proposed here are closely related to three different approaches for the solution of inverse problems. First, several studies have incorporated diffusion models as regularizers in inverse problems [yang2022diffusion, darassurvey]. Chung et al. [chung2022diffusion] proposed diffusion posterior sampling strategy to efficiently solve noisy non-linear inverse problems via approximation of the posterior sampling. Chung et al. [chung2022come] introduced Come-Closer-Diffuse-Faster technique to show how the feed-forward neural networks for inverse problems can be synergistically combined with the diffusion models for a variety of problems like image inpainting and super-resolution. Song et al. [song2022solving] proposed a fully unsupervised method for inverse problems based on score-matching diffusion models, which reconstructs an image consistent with both the prior and the observed measurements. Li et al. [li2022srdiff] proposed SRDiff, a diffusion-based model that generates high-resolution images from low-resolution inputs by formulating single image super-resolution as a conditional diffusion process. Whang et al. [whang2022deblurring] introduced a technique for deblurring using conditional diffusion models which trains a stochastic sampler that refined the output of a deterministic predictor, producing a diverse set of plausible reconstructions for a given input, thereby improving image perceptual quality. The key idea often used in these methods was to leverage a pretrained diffusion model that captures the data distribution as a prior, and then condition this model on the given measurements to infer the underlying clean signal or image. For inverse problems, diffusion methods condition the diffusion model on the given measurements (e.g. noisy, incomplete, or compressed data) by incorporating them into the denoising process. Nonetheless, it has been shown in [eliasof2024completedeeplearningmethod] that pretrained diffusion models that are used for ill-posed inverse problems as regularizers tend to under-perform as compared to the models that are trained specifically on a particular inverse problem. In particular, such models tend to break when the noise level is not very low.
A second branch of techniques that are related to the work proposed here uses encoder–decoder style neural architectures for solving inverse problems. For instance, Chung et al. [chung2024pairedautoencodersinverseproblems] proposed paired-autoencoders for solving inverse problems that learns joint latent representations of data and model spaces, enabling likelihood-free estimation for PDE-based inverse problems while also providing refinement strategies. Radev et al. [radev2020bayesflow] proposed an invertible neural network combined with a learned summary (encoder) network to map observations to latent parameters, enabling amortized posterior inference without explicit likelihoods. Wang et al. [wang2021theory] proposed theory-guided autoencoder which discretized PDE residuals into an autoencoder loss, enabling surrogate modeling and inverse parameter estimation for subsurface flow with limited labeled data. Padmanabha et al. [padmanabha2021solving] constructed a conditional invertible neural network that maps measurements to latent variables (and vice versa) under a learned bijective transformation, allowing likelihood-free inversion by running the invertible mapping backward to recover parameter estimates conditioned on observations. Together, these works illustrate how encoder–decoder architectures can be harnessed as powerful likelihood-free estimators, enabling direct data-to-solution mappings that bypass explicit likelihood evaluation while tackling complex inverse problems. This property is especially advantageous when the forward operator is either computationally intractable or analytically unavailable, which is often the case in high-dimensional imaging applications.
A third branch of techniques that relates to our approach integrates the forward problem directly into the neural network architecture. In this line of work, the forward operator is embedded into the training loop so that the data misfit (and its gradient) can be computed within the network and used as an explicit guide for learning. For instance, Eliasof et al. [eliasof2024completedeeplearningmethod] introduced InverseUNetODE, which incorporates the adjoint of the forward model at each layer of a UNet to iteratively refine reconstructions, demonstrating improved stability in highly ill-posed settings. Mukherjee et al. [mukherjee2021learning] proposed learning convex regularizers that satisfy variational source conditions, combining deep learning with principles of classical variational analysis to enforce forward-consistency during inversion. Similarly, Eliasof et al. [eliasof2023drip] introduced a method called DRiP, a variational framework that learns neural regularization functionals guaranteeing data consistency by combining least-action principles with deep neural architectures, thereby overcoming the instability of proximal-based methods. Jin et al. [jin2017deep] introduced FBPConvNet, which combines filtered back projection with a U-Net-style CNN to remove artifacts from ill-posed inverse reconstructions, effectively embedding the forward operator into the learning process to ensure physics-consistent solutions. Collectively, these methods highlight the value of embedding the forward model within the network, as they not only reduce the reliance on external priors but also ensure that learned reconstructions remain anchored to the measured data. Our approach utilizes components from this methodology by incorporating forward-model–based misfit terms into the training of flow-matching networks, thereby grounding the generative process in the measured observations.
Additionally, a complementary line of research has recently focused on guided flow matching techniques, which extend the basic FM formalism by introducing additional control signals or constraints during training and sampling. Zheng et al. [zheng2023guided] proposed Guided Flows, incorporating classifier-free guidance into velocity fields to achieve conditional generative modeling across images, speech, and decision-making tasks. Chemseddine et al. [chemseddine2025conditional] introduced conditional Wasserstein distance that aligns posterior distributions by forbidding mass transport in the observation space, thereby enabling Bayesian optimal transport-based flows. Ben-Hamu et al. [ben2024d] introduced D-Flow, a framework that differentiates through the diffusion/FM flow to enable controlled generation by backpropagating task-specific objectives D-Flow, a framework that controls generation in diffusion and flow-matching models by differentiating through the flow, optimizing for source point. Furthermore, Pokle et al. [pokle2023training] demonstrated training-free linear image inversion using pretrained flow models, where gradient correction techniques and conditional optimal transport probability paths are used to solve inverse problems. Together, these works illustrate the versatility of guided flow matching, highlighting how auxiliary signals - ranging from classifier-based gradients to transport constraints - can steer the generative dynamics toward desired data distributions or application-specific goals. Similarly, our work also introduces guidance via incorporating measured data as well as noise variance directly in the network architecture and optimizing it using measured data misfit to direct the flow toward solving inverse problems in imaging.
@@
Main contribution: The core contributions of this paper lie in the design and application of the Data-AWare and Noise-Informed Flow Matching (DAWN-FM) framework, specifically crafted for solving highly ill-posed inverse problems like image deblurring and tomography. (i) Our method is designed to be problem-specific, which adjusts itself to the unique structure of the inverse problem it is tasked with, learning the posterior distribution directly, ensuring that the learned velocity fields and mappings are directly applicable to the target task. (ii) We train an interpolant by embedding measured data and noise information directly into the interpolation process. This incorporation allows our model to adapt to a wide range of noise conditions, something that pretrained models struggle with. By making the training explicitly aware of the noise level and data characteristics, our model can better navigate noisy or incomplete measured data, producing superior reconstructions. (iii) By learning the posterior distribution directly and leveraging the stochastic nature of the interpolation process, DAWN-FM generates multiple plausible solutions for a given inverse problem, allowing us to explore the solution space more thoroughly and in particular, to estimate the posterior mean and its standard deviation, estimating the uncertainty in the recovered solution.
Organization of the paper: The remainder of this paper is organized as follows. In Section 2, we review the fundamentals of Flow Matching and describe how it can be adapted to solve inverse problems, including the design of our proposed data-aware and noise-informed velocity estimator. Section 3 provides mathematical details for the two inverse problems we solve in this work – image deblurring and tomography. Section 4 presents numerical experiments on image deblurring and tomography, evaluating the performance of the proposed methods against existing approaches. Section 5 concludes the paper with a summary of findings and a discussion on future research directions. Additional technical details, including evaluation metrics, experimental settings and additional visualizations are provided in the Supplementary Material (Sections 6, 7, 8 and 9).
2. Flow Matching and Inverse Problems
@@
2.1. Flow Matching (FM): A partial review
Flow Matching (FM) is a framework that transforms points between two distributions. Given two densities $\pi_{0}({\bf x})$ and $\pi_{1}({\bf x})$, the goal is to find a mapping that takes a point ${\bf x}_{0}\sim\pi_{0}({\bf x})$ and transports it to a point ${\bf x}_{1}\sim\pi_{1}({\bf x})$. For simplicity and for the purpose of this work, we choose $\pi_{0}$ to be a Gaussian distribution with 0 mean and $\bf I$} covariance.
Following the linear interpolation described in [liu2022flow], we define trajectories parametrized by the time variable $t$ as,
${\bf x}_{t}=(1-t){\bf x}_{0}+t{\bf x}_{1},$ (1)
where ${\bf x}_{t}\sim\pi_{t}({\bf x})$ is the linear interpolation of ${\bf x}_{0}$ and ${\bf x}_{1}$ at any given time $t$. These trajectories connect points from ${\bf x}_{0}$ at $t=0$ to ${\bf x}_{1}$ at $t=1$, i.e., the parameter $t\in[0,1]$ interpolates between the reference distribution and target distribution and appears explicitly in the velocity estimator (see Equation 4) to ensure that the learned flow evolves smoothly and consistently along this continuum. More complex trajectories have been proposed in [albergo2023stochastic], however, in our context, we found that simple linear trajectories suffice and have advantages for being very smooth in time. The velocity along the trajectory is the time derivative of the interpolant ${\bf x}_{t}$, that is,
${\frac{d{\bf x}_{t}}{dt}}={\bf v}={\bf x}_{1}-{\bf x}_{0}.$ (2)
Using the velocity vector defined in Equation 2, we aim drive the flow to follow the direction $({\bf x}_{1}-{\bf x}_{0})$ of the linear path pointing from ${\bf x}_{0}\sim\pi_{0}$ to ${\bf x}_{1}\sim\pi_{1}$. To this end, we parameterize the velocity by a learnable function ${\bf s}_{{\boldsymbol{\theta}}}({\bf x}_{t},t)$ and solve the optimization problem for $\boldsymbol{\theta}$} as,
$\hat{\boldsymbol{\theta}}={\rm arg}\min_{{\boldsymbol{\theta}}}{\mathbb{E}}_{{\bf x}_{0},{\bf x}_{1},t}\Big[\|{\bf s}_{\theta}({\bf x}_{t},t)-{\bf v}\|^{2}\Big]={\rm arg}\min_{{\boldsymbol{\theta}}}{\mathbb{E}}_{{\bf x}_{0},{\bf x}_{1},t}\Big[\|{\bf s}_{\theta}({\bf x}_{t},t)+{\bf x}_{0}-{\bf x}_{1}\|^{2}\Big],$ (3)
where $t\sim\mathcal{U}[0,1]$ is a uniform random variable [lipman2022flow, liu2022flow]. After training, the velocity function ${\bf s}_{{\boldsymbol{\theta}}}({\bf x}_{t},t)$ can be estimated for every point in time. In this work, given ${\bf x}_{0}$, we recover ${\bf x}_{1}$ by numerically integrating the ODE
@@ 2.1. Flow Matching (FM): A partial review
$\hat{{\bf x}}_{1}={\bf x}_{t}+(1-t){\bf v}\approx{\bf x}_{t}+(1-t){\bf s}_{{\boldsymbol{\theta}}}({\bf x}_{t},t).$ (5)
@@
2.2. Applying FM for solving inverse problems
@@
${\bf A}{\bf x}_{1}+{\boldsymbol{\varepsilon}}={\bf b}.$ (6)
Here, $\bf A$} is a linear forward mapping (although the method can work for nonlinear mappings as well) and ${\boldsymbol{\varepsilon}}\sim\mathcal{N}(0,\sigma^{2}{\bf I})$ is a random vector. We assume that $\bf A$} is rank-deficient or numerically rank-deficient [hansen], so the effective dimension of $\bf b$} is smaller than ${\bf x}_{1}$ and one cannot obtain a reasonable estimate for ${\bf x}_{1}$ given the noisy data $\bf b$} without the incorporation of a-priori information.
@@
Bayes’ theorem suggests that it is possible to factor the posterior distribution $\pi({\bf x}_{1}|{\bf b})$ using the known distribution of $\pi({\bf b}|{\bf x}_{1})$ and the prior distribution $\pi({\bf x}_{1})$. This observation motivated a number of studies that used the estimated pretrained distribution $\pi({\bf x}_{1})$ in the process of solving an inverse problem [yang2022diffusion, chung2022diffusion, chung2022come, chung2022improving, song2022solving]. Nonetheless, it has been shown in [eliasof2024completedeeplearningmethod] that such estimators tend to produce unsatisfactory results when solving highly ill-posed problems or when the data is very noisy. The reason for this behaviour stems from the fact that pretrained estimators push the solution towards the center of the prior distribution, irrespective of what the data represents. To see this, we use a careful dissection of the solution. Consider the singular value decomposition
${\bf A}=\sum_{i}{\boldsymbol{\lambda}}_{i}{\bf u}_{i}{\bf v}_{i}^{\top},$
@@
where ${\bf u}_{i}$ and ${\bf v}_{i}$ are the left and right singular vectors and ${\boldsymbol{\lambda}}_{i}$ are the singular values. Given the orthogonality of ${\bf u}_{i}$, we can decouple the data equations into
${\boldsymbol{\lambda}}_{i}({\bf v}_{i}^{\top}{\bf x}_{1})=({\bf u}_{i}^{\top}{\bf b}),\quad i=1,\ldots,n.$
@@
If ${\boldsymbol{\lambda}}_{i}$ is large, the projection of ${\bf x}_{1}$ onto the eigenvector ${\bf v}_{i}$ is very informative and very minimal regularization is required. However, if ${\boldsymbol{\lambda}}_{i}\approx 0$, the contribution of ${\bf v}_{i}$ to the solution is difficult, if not impossible, to obtain and this is where the regularization is highly needed. When using pretrained models, the prior $\pi({\bf x}_{1})$ is estimated numerically and it is unaware of the inverse problem at hand. Errors in the estimated $\pi({\bf x}_{1})$ in the parts that correspond to the large singular vectors may not be destructive. However, if $\pi({\bf x}_{1})$ has errors that correspond to the very small singular values, that is, to the effective null space of the data, this may lead to the artifacts that have been observed early in [somersallo, Tenorio2011]. This suggests that, although it is appealing to use a generic pretrained priors in the process, a better approach is to not use the Bayesian factorization to prior and likelihood but rather train a flow matching interpolant that maps the distribution $\pi({\bf x}_{0})$ to the posterior $\pi({\bf x}_{1}|{\bf b})$ directly. Indeed, as we show in Section 2.3, the flexibility of the FM framework allows us to learn a velocity function that achieves just that.
Inverse crime considerations. In the original formulation of our inverse problem setup, the measured data $\bf b$} is generated via ${\bf b}={\bf A}{\bf x}_{1}+{\boldsymbol{\varepsilon}}$, where $\bf A$} is the same forward operator subsequently used during inversion. This formulation, while standard in many machine learning-based inverse problem studies, can inadvertently lead to an inverse crime [wirgin2004inverse] - a situation where the forward model used for synthetic data generation is identical to that used for inversion, potentially yielding overly optimistic reconstruction results. To rigorously assess the robustness of our proposed DAWN-FM framework under more realistic conditions, we performed additional experiments (reported in Section 7 under Supplementary Material) in which the forward model used for data generation was intentionally modified. Specifically, we employed a composite operator $\tilde{{\bf A}}={\bf D}{\bf A}{\bf U}$ where $\bf D$} denotes a downsampling operator, $\bf U$} an upsampling operator, with $\bf A$} being the original forward model. This change ensures that the forward mapping used to synthesize the training and test data differs from that used during inversion, thereby eliminating the risk of inverse crime. Our results confirm that the effectiveness of our method does not depend on any implicit inverse crime, and the method generalizes well even when the forward model used for data generation is perturbed. For detailed discussion and results considering the issue of inverse crime, refer to Table 3 under Section 7 in the Supplementary Material.
2.3. A data-aware and noise-informed velocity estimator
Figure 1. Schematics for standard flow matching (FM) (left) and flow matching for solving inverse problem characterized by the forward problem $\bf A$} along with an additive noise scale $\sigma$ (right). Here, ${\bf s}_{\boldsymbol{\theta}}$ represents the trained network for estimating the velocity along the trajectory ${\bf x}_{t}$ at time $t$ between the reference distribution $\pi_{0}$ and the target distribution $\pi_{1}$, and $f$ represents a transformation on the measured (noisy) data which is used as data-embedding into the flow-matching network on the right.